OpenStudy (anonymous):

find he discriminant of 2x^2+3x=0

OpenStudy (anonymous):

\(2x^2+3x=0\) \(ax^2+bx+c=0\) \(a=2,b=3,c=0\) compute \(b^2-4ac\) get \(3^2=9\)